OpenStudy (anonymous):

x + y = k x - y = k The solution of the system shown is _____. (0, k) (0, 0) (k, 0)

OpenStudy (anonymous):

im not sure it's correct, but by using solving system of equations you should get (k, 0)

OpenStudy (maheshmeghwal9):

(k,o) may be the right answer.

OpenStudy (maheshmeghwal9):

Is that fine to you?

OpenStudy (maheshmeghwal9):

Reply hithere!

OpenStudy (zepp):

Let's use the reduction method by adding them up, you'll cancel out the y's, leaving you with x=k Now if x=k, let's replace k by x x-y=k x-y=x x-x=y 0=y Since x = k and 0 = y, we'll say that the solution to this system is (K,0)